\begin{thebibliography}{99}                                                                                               %
\bibitem {MR1972755}M.~Aunola. \newblock The discretized harmonic oscillator:
{M}athieu functions and a new  class of generalized {H}ermite polynomials.
\newblock {\em J. Math. Phys.}, 44(5):1913--1936, 2003.

\bibitem {charlier}D.~Dominici. \newblock Asymptotic analysis of the
{A}skey-scheme {I}: from {K}rawtchouk to  {C}harlier. \newblock Preprint,
arXiv: math.CA/0501072, 2005.

\bibitem {MR0058756}A.~Erd{\'e}lyi, W.~Magnus, F.~Oberhettinger, and F.~G.
Tricomi. \newblock {\em Higher transcendental functions. {V}ols. {I}, {II}}.
\newblock McGraw-Hill Book Company, Inc., New York-Toronto-London, 1953.

\bibitem {MR1985821}C.~Ferreira, J.~L. Lopez, and E.~Mainar.
\newblock Asymptotic relations in the {A}skey scheme for hypergeometric
orthogonal polynomials. \newblock {\em Adv. in Appl. Math.}, 31(1):61--85, 2003.

\bibitem {MR2076056}A.~Gil, J.~Segura, and N.~M. Temme. \newblock Integral
representations for computing real parabolic cylinder  functions.
\newblock {\em Numer. Math.}, 98(1):105--134, 2004.

\bibitem {hermpoly}C.~Hermite. \newblock Sur un nouveau d\'eveloppement en
s\'erie de functions. \newblock {\em Compt. Rend. Acad. Sci. Paris},
58:93--100, 1864.

\bibitem {MR1736478}R.~Hinterding and J.~Wess. \newblock {$q$}-deformed
{H}ermite polynomials in {$q$}-quantum mechanics.
\newblock {\em Eur. Phys. J. C Part. Fields}, 6(1):183--186, 1999.

\bibitem {koekoek94askeyscheme}R.~Koekoek and R.~F. Swarttouw. \newblock The
{A}skey-scheme of hypergeometric orthogonal polynomials and its  $q$-analogue.
\newblock Technical Report 98-17, Delft University of Technology, 1998. \newblock http://aw.twi.tudelft.nl/~koekoek/askey/.

\bibitem {MR0350075}N.~N. Lebedev.
\newblock {\em Special functions and their applications}. \newblock Dover
Publications Inc., New York, 1972.

\bibitem {MR1803886}J.~L. L{\'o}pez and N.~M. Temme. \newblock Approximation
of orthogonal polynomials in terms of {H}ermite  polynomials.
\newblock {\em Methods Appl. Anal.}, 6(2):131--146, 1999. \newblock Dedicated
to Richard A.\ Askey on the occasion of his 65th birthday,  Part II.

\bibitem {MR1723073}J.~L. L{\'o}pez and N.~M. Temme. \newblock Hermite
polynomials in asymptotic representations of generalized  {B}ernoulli,
{E}uler, {B}essel, and {B}uchholz polynomials.
\newblock {\em J. Math. Anal. Appl.}, 239(2):457--477, 1999.

\bibitem {MR2068115}J.~L. L{\'o}pez and N.~M. Temme. \newblock Convergent
asymptotic expansions of {C}harlier, {L}aguerre and  {J}acobi polynomials.
\newblock {\em Proc. Roy. Soc. Edinburgh Sect. A}, 134(3):537--555, 2004.

\bibitem {MR810399}L.~Pauling and E.~B. Wilson, Jr.
\newblock {\em Introduction to quantum mechanics}. \newblock Dover
Publications Inc., New York, 1985.

\bibitem {atlas}J.~Spanier and K.~B. Oldham.
\newblock {\em An Atlas of Functions}. \newblock Hemisphere Pub. Corp., 1987.

\bibitem {MR0372517}G.~Szeg{\H{o}}. \newblock {\em Orthogonal polynomials}.
\newblock American Mathematical Society, Providence, R.I., fourth edition,  1975.

\bibitem {MR1036513}N.~M. Temme. \newblock Asymptotic estimates for {L}aguerre
polynomials. \newblock {\em Z. Angew. Math. Phys.}, 41(1):114--126, 1990.

\bibitem {MR1052447}N.~M. Temme. \newblock Polynomial asymptotic estimates of
{G}egenbauer, {L}aguerre, and  {J}acobi polynomials. \newblock In
\emph{Asymptotic and computational analysis (Winnipeg, MB, 1989)},  volume 124
of \emph{Lecture Notes in Pure and Appl. Math.}, pages 455--476.  Dekker, New
York, 1990.

\bibitem {MR1376370}N.~M. Temme. \newblock {\em Special functions}.
\newblock A Wiley-Interscience Publication. John Wiley \& Sons Inc., New
York,  1996.

\bibitem {MR1780050}N.~M. Temme. \newblock Numerical and asymptotic aspects of
parabolic cylinder functions. \newblock {\em J. Comput. Appl. Math.},
121(1-2):221--246, 2000. \newblock Numerical analysis in the 20th century,
Vol.\ I, Approximation  theory.

\bibitem {MR1805994}N.~M. Temme and J.~L. L{\'o}pez. \newblock The role of
{H}ermite polynomials in asymptotic analysis. \newblock In \emph{Special
functions (Hong Kong, 1999)}, pages 339--350. World  Sci. Publishing, River
Edge, NJ, 2000.

\bibitem {MR1858318}N.~M. Temme and J.~L. L{\'o}pez. \newblock The {A}skey
scheme for hypergeometric orthogonal polynomials viewed  from asymptotic
analysis. \newblock {\em J. Comput. Appl. Math.}, 133(1-2):623--633, 2001.
\newblock Proceedings of the Fifth International Symposium on Orthogonal
Polynomials, Special Functions and their Applications (Patras, 1999).

\bibitem {MR1993339}N.~M. Temme and R.~Vidunas. \newblock Parabolic cylinder
functions: examples of error bounds for asymptotic  expansions.
\newblock {\em Anal. Appl. (Singap.)}, 1(3):265--288, 2003.

\bibitem {MR1349110}G.~N. Watson.
\newblock {\em A treatise on the theory of {B}essel functions}.
\newblock Cambridge Mathematical Library. Cambridge University Press,
Cambridge, 1995.

\bibitem {MR1343548}A.~Zarzo, J.~S. Dehesa, and J.~Torres. \newblock On a new
set of polynomials representing the wave functions of the  quantum
relativistic harmonic oscillator. \newblock {\em Ann. Numer. Math.},
2(1-4):439--455, 1995. \newblock Special functions (Torino, 1993).
\end{thebibliography}

\documentclass[12pt]{article}%
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{amssymb}%
\newtheorem{theorem}{Theorem}

\begin{document}
\title{Asymptotic analysis of the Askey-scheme II: from Charlier to Hermite}
\author{Diego Dominici \thanks{e-mail: dominicd@newpaltz.edu}\\Department of Mathematics\\State University of New York at New Paltz\\75 S. Manheim Blvd. Suite 9\\New Paltz, NY 12561-2443\\USA\\Phone: (845) 257-2607\\
Fax: (845) 257-3571
}

\maketitle

\begin{abstract}
We analyze the Hermite polynomials $H_{n}(\xi)$ and their zeros asymptotically
as $n\rightarrow\infty,$ using the limit relation between the Charlier and
Hermite polynomials. Our formulas involve some special functions and they
yield very accurate approximations.
\end{abstract}

Keywords: Hermite polynomials, Askey-scheme, asymptotic analysis, orthogonal
polynomials, hypergeometric polynomials, special functions.

MSC-class: 33C45 (Primary) 34E05, 33C10 (Secondary)


\section{Introduction}

The Hermite polynomials $H_{n}(x)$ are defined by \cite{MR1376370}
\begin{equation}
H_{n}(x)=n!%
{\displaystyle\sum\limits_{k=0}^{\left\lfloor \frac{n}{2}\right\rfloor }}
\frac{\left(  -1\right)  ^{k}}{k!(n-2k)!}\left(  2x\right)  ^{n-2k}
\label{Cndef}%
\end{equation}
for $n=0,1,\ldots.$ They satisfy the orthogonality condition \cite{atlas}
\[%
{\displaystyle\int\limits_{-\infty}^{\infty}}
e^{-x^{2}}H_{m}(x)H_{n}(x)dx=\sqrt{\pi}2^{n}n!\delta_{nm}%
\]
and the reflection formula
\begin{equation}
H_{n}(-x)=(-1)^{n}H_{n}(x). \label{reflec}%
\end{equation}
The Hermite polynomials are special cases of the parabolic cylinder function
$U(a,z),$
\[
H_{n}(x)=2^{\frac{n}{2}}\exp\left(  \frac{x^{2}}{2}\right)  U\left(
-n-\frac{1}{2},\sqrt{2}x\right) ,
\]
which was analyzed by Nico Temme in \cite{MR2076056}, \cite{MR1780050} and
\cite{MR1993339}.

The Hermite polynomials have been extensively studied since the pioneer
article of C. Hermite \cite{hermpoly} in 1864 (they were previously considered
by Fourier and Chebyshev). They have many applications in the physical
sciences and are particularly important in the quantum mechanical treatment of
the harmonic oscillator \cite{MR810399} (see also \cite{MR1972755},
\cite{MR1736478} and \cite{MR1343548} for some extensions). We refer the
interested reader to \cite{MR0058756} and \cite{MR0372517} for further
properties and references.

There are several families of orthogonal polynomials which have asymptotic
approximations in terms of $H_{n}(x).$ Some cases studied by Nico Temme
include the Gegenbauer \cite{MR1052447}, Laguerre \cite{MR2068115},
\cite{MR1036513}, Tricomi-Carlitz and Jacobi polynomials \cite{MR1803886}. He
also considered the asymptotic representations of other families of
polynomials such as the generalized Bernoulli, Euler, Bessel and Buchholz
polynomials in \cite{MR1723073}.

A rich source of asymptotic relations between the $H_{n}(x)$ and other
polynomials \cite{MR1985821}, \cite{MR1805994}, \cite{MR1858318} is provided
by the Askey-scheme of hypergeometric orthogonal polynomials
\cite{koekoek94askeyscheme}:

$%
\begin{array}
[c]{ccccccccc}%
_{4}F_{3} & \fbox{Wilson} & \  & \fbox{\ Racah} &  &  &  &  & \\
& \downarrow\quad\searrow &  & \downarrow\quad\searrow &  &  &  &  & \\
_{3}F_{2} & \ \fbox{$%
\begin{array}
[c]{c}%
\text{Continuous }\\
\text{dual Hahn}%
\end{array}
$} & \fbox{$%
\begin{array}
[c]{c}%
\text{Continuous }\\
\text{Hahn}%
\end{array}
$}\  & \fbox{Hahn} & \fbox{Dual Hahn} &  &  &  & \\
& \downarrow & \swarrow\quad\downarrow & \swarrow\quad\downarrow\quad\searrow
& \swarrow\quad\downarrow &  &  &  & \\
_{2}F_{1} & \fbox{$%
\begin{array}
[c]{c}%
\text{Meixner}\\
\text{Pollaczek}%
\end{array}
$} & \fbox{Jacobi} & \ \fbox{Meixner} & \fbox{Krawtchouk} &  &  &  & \\
& \searrow & \downarrow & \swarrow\quad\searrow & \downarrow &  &  &  & \\
_{1}F_{1} &  & \fbox{Laguerre} &  & \fbox{Charlier\ } & _{2}F_{0} &  &  & \\
&  & \searrow &  & \swarrow &  &  &  & \\
_{2}F_{0} &  &  & \fbox{Hermite} &  &  &  &  &
\end{array}
$

where the arrows indicate limit relations between the polynomials.

In particular, the limit relation between the Charlier polynomials
$C_{n}^{(a)}(x)$ and the Hermite polynomials is given by%
\begin{equation}
\underset{a\rightarrow\infty}{\lim}\left(  -1\right)  ^{n}\left(  2a\right)
^{\frac{n}{2}}C_{n}^{(a)}\left(  a+x\sqrt{2a}\right)  =H_{n}(x).\label{limit}%
\end{equation}

In this article we investigate the asymptotic behavior of $H_{n}(x)$ as
$n\rightarrow\infty,$ using (\ref{limit}) and the asymptotic results on the
Charlier polynomials derived in \cite{charlier}. We believe that our method
provides a useful approach to asymptotic analysis and could be used for other
families of polynomials of the Askey scheme.

\section{Previous results}

We define the Charlier polynomials by
\begin{equation}
\,C_{n}^{(a)}(x)=\,_{2}F_{0}\left(  \left.
\begin{array}
[c]{c}%
-n,-x\\
-
\end{array}
\right\vert -\frac{1}{a}\right)  ,\quad n=0,1,\ldots\label{Charlier}%
\end{equation}
with $a>0.$ The following results were derived in \cite{charlier}.

\begin{theorem}
As $n\rightarrow\infty,$ $C_{n}^{(a)}(x)$ admits the following asymptotic
approximations, with%
\begin{equation}
\Omega^{\pm}=\left(  \sqrt{n}\pm\sqrt{a}\right)  ^{2}. \label{Omega}%
\end{equation}

\begin{enumerate}
\item $n=O(1).$%
\begin{equation}
C_{n}^{(a)}\simeq\left(  1-\frac{x}{a}\right)  ^{n} \label{c1}%
\end{equation}

\item $x<\Omega^{-},\quad0<n<a.$
\begin{equation}
C_{n}^{(a)}\sim F_{3}(x)=\exp\left[  \Psi_{3}(x)\right]  L_{3}(x), \label{c2}%
\end{equation}
where
\begin{equation}
\Psi_{3}(x)=x\ln\left(  \frac{a+x-n+\Delta}{2a}\right)  +n\ln\left(
\frac{a-x+n+\Delta}{2a}\right)  +\frac{1}{2}\left(  a-x-n+\Delta\right)
\label{psi3}%
\end{equation}
and%
\begin{equation}
L_{3}(x)=\sqrt{\frac{a-x-n+\Delta}{2\Delta}}, \label{l3}%
\end{equation}
with%
\begin{equation}
\Delta(x)=\sqrt{a^{2}-2a(x+n)+(x-n)^{2}}. \label{delta}%
\end{equation}

\item $\Omega^{+}<x.$
\begin{equation}
C_{n}^{(a)}\sim F_{4}(x)=\left(  -1\right)  ^{n}\exp\left[  \Psi
_{4}(x)\right]  L_{4}(x), \label{c4}%
\end{equation}
where%
\begin{equation}
\Psi_{4}(x)=x\ln\left(  \frac{a+x-n-\Delta}{2a}\right)  +n\ln\left(
\frac{x-a-n+\Delta}{2a}\right)  +\frac{1}{2}\left(  a-x-n+\Delta\right)
\label{psi4}%
\end{equation}
and%
\begin{equation}
L_{4}(x)=\sqrt{\frac{x-a+n+\Delta}{2\Delta}}. \label{l4}%
\end{equation}

\item $x\approx\Omega^{-},\ 0<n<a.$%
\begin{align}
C_{n}^{(a)}  &  \sim\sqrt{2\pi}\left(  \frac{n}{a}\right)  ^{\frac{1}{6}%
}\left(  \sqrt{a}-\sqrt{n}\right)  ^{\frac{1}{3}}\mathrm{Ai}\left[  \left(
\frac{n}{a}\right)  ^{\frac{1}{6}}\frac{\left(  \Omega^{-}-x\right)  }{\left(
\sqrt{a}-\sqrt{n}\right)  ^{\frac{2}{3}}}\right] \label{c8}\\
&  \times\exp\left[  \frac{1}{2}n\ln\left(  \frac{n}{a}\right)  +x\ln\left(
1-\sqrt{\frac{n}{a}}\right)  +\sqrt{an}-\sqrt{n}\right]  ,\nonumber
\end{align}
where $\mathrm{Ai}\left(  \cdot\right)  $ is the Airy function.

\item $\Omega^{-}<x<\Omega^{+}.$%
\begin{equation}
C_{n}^{(a)}\sim F_{10}(x)=F_{3}(x)+F_{4}(x). \label{c10}%
\end{equation}

\item $x\approx\Omega^{+}.$%
\begin{align}
C_{n}^{(a)}  &  \sim\sqrt{2\pi}\left(  \frac{n}{a}\right)  ^{\frac{1}{6}%
}\left(  \sqrt{a}+\sqrt{n}\right)  ^{\frac{1}{3}}\mathrm{Ai}\left[  \left(
\frac{n}{a}\right)  ^{\frac{1}{6}}\frac{\left(  x-\Omega^{+}\right)  }{\left(
\sqrt{a}+\sqrt{n}\right)  ^{\frac{2}{3}}}\right] \label{c11}\\
&  \times\left(  -1\right)  ^{n}\exp\left[  \frac{1}{2}n\ln\left(  \frac{n}%
{a}\right)  +x\ln\left(  1+\sqrt{\frac{n}{a}}\right)  -\sqrt{an}-\sqrt
{n}\right]  .\nonumber
\end{align}

\end{enumerate}
\end{theorem}

\section{Limit analysis}

From (\ref{Omega}) we have, as $a\rightarrow\infty,$
\begin{equation}
\frac{\Omega^{\pm}-a}{\sqrt{2a}}\rightarrow\pm\sqrt{2n}. \label{limitOmega}%
\end{equation}
Hence, the six regions of Theorem 1 transform into the following regions:

\begin{enumerate}
\item Region I: $n=O(1).$

Setting
\begin{equation}
x=a+\xi\sqrt{2a} \label{eta}%
\end{equation}
in (\ref{c1}) and using (\ref{limit}), we get%
\begin{equation}
H_{n}(\xi)\simeq\left(  2\xi\right)  ^{n}. \label{h1}%
\end{equation}
The formula above is exact for $n=0,1$ and is a good approximation when
$\xi\gg n.$

\item Region II: $\xi<-\sqrt{2n}.$

From (\ref{delta}) and (\ref{eta}) we have
\begin{equation}
\Delta\sim\sqrt{2}\sigma\sqrt{a}-\frac{\xi n}{\sigma},\quad a\rightarrow
\infty, \label{delta1}%
\end{equation}
with%
\begin{equation}
\sigma=\sqrt{\xi^{2}-2n}. \label{sigma}%
\end{equation}
Using (\ref{delta1}) in (\ref{psi3}) and (\ref{l3}), we get%
\begin{equation}
\frac{n}{2}\ln(2a)+\Psi_{3}(x)\sim\Phi_{1}\left(  \xi\right)  \equiv
n\ln\left(  \sigma-\xi\right)  +\frac{\xi^{2}+\sigma\xi-n}{2}, \label{psi3lim}%
\end{equation}%
\begin{equation}
L_{3}(x)\sim U_{1}\left(  \xi\right)  =\sqrt{\frac{1}{2}\left(  1-\frac{\xi
}{\sigma}\right)  ,} \label{l3lim}%
\end{equation}
as $a\rightarrow\infty.$ Thus, from (\ref{limit}) we have%
\begin{equation}
H_{n}(\xi)\sim\Lambda_{1}\left(  \xi\right)  \equiv\left(  -1\right)  ^{n}%
\exp\left[  \Phi_{1}\left(  \xi\right)  \right]  U_{1}\left(  \xi\right)  .
\label{H2}%
\end{equation}

\item Region III: $\xi>\sqrt{2n}.$

Using (\ref{delta1}) in (\ref{psi4}) and (\ref{l4}), we obtain%
\begin{equation}
\frac{n}{2}\ln(2a)+\Psi_{4}(x)\sim\Phi_{2}\left(  \xi\right)  \equiv
n\ln\left(  \sigma+\xi\right)  +\frac{\xi^{2}-\sigma\xi-n}{2}, \label{psi4lim}%
\end{equation}%
\begin{equation}
L_{4}(x)\sim U_{2}\left(  \xi\right)  =\sqrt{\frac{1}{2}\left(  1+\frac{\xi
}{\sigma}\right)  ,} \label{l4lim}%
\end{equation}
as $a\rightarrow\infty.$ Hence,%
\begin{equation}
H_{n}(\xi)\sim\Lambda_{2}\left(  \xi\right)  \equiv\exp\left[  \Phi_{2}\left(
\xi\right)  \right]  U_{2}\left(  \xi\right)  . \label{H3}%
\end{equation}
Note that
\[
\Lambda_{1}\left(  -\xi\right)  =\left(  -1\right)  ^{n}\Lambda_{2}\left(
\xi\right)  ,
\]
as one would expect from (\ref{reflec}).

\item Region IV: $\xi\approx-\sqrt{2n}.$

Using (\ref{eta}) in (\ref{c8}) we have, as $a\rightarrow\infty,$%
\[
\frac{n}{2}\ln(2a)+\frac{1}{2}n\ln\left(  \frac{n}{a}\right)  +x\ln\left(
1-\sqrt{\frac{n}{a}}\right)  +\sqrt{an}-\sqrt{n}\sim\Phi_{3}\left(
\xi\right)  ,
\]
where%
\begin{equation}
\Phi_{3}\left(  \xi\right)  =\frac{n}{2}\ln\left(  2n\right)  -\frac{3}%
{2}n-\xi\sqrt{2n}. \label{phi3}%
\end{equation}
Also,%
\[
\sqrt{2\pi}\left(  \frac{n}{a}\right)  ^{\frac{1}{6}}\left(  \sqrt{a}-\sqrt
{n}\right)  ^{\frac{1}{3}}\sim n^{\frac{1}{6}}%
\]
and%
\[
\left(  \frac{n}{a}\right)  ^{\frac{1}{6}}\frac{\left(  \Omega^{-}-x\right)
}{\left(  \sqrt{a}-\sqrt{n}\right)  ^{\frac{2}{3}}}\sim-n^{\frac{1}{6}}%
\sqrt{2}\left(  \xi+\sqrt{2n}\right)  .
\]
Therefore,%
\begin{equation}
H_{n}(\xi)\sim\Lambda_{3}\left(  \xi\right)  \equiv(-1)^{n}\sqrt{2\pi}%
n^{\frac{1}{6}}\exp\left[  \Phi_{3}\left(  \xi\right)  \right]  \mathrm{Ai}%
\left[  -n^{\frac{1}{6}}\sqrt{2}\left(  \xi+\sqrt{2n}\right)  \right]  .
\label{H4}%
\end{equation}

\item Region V: $\xi\approx\sqrt{2n}.$

Using (\ref{eta}) in (\ref{c11}) we have, as $a\rightarrow\infty,$%
\[
\frac{n}{2}\ln(2a)+\frac{1}{2}n\ln\left(  \frac{n}{a}\right)  +x\ln\left(
1+\sqrt{\frac{n}{a}}\right)  -\sqrt{an}-\sqrt{n}\sim\Phi_{4}\left(
\xi\right)  ,
\]
with%
\begin{equation}
\Phi_{4}\left(  \xi\right)  =\frac{n}{2}\ln\left(  2n\right)  -\frac{3}%
{2}n-\xi\sqrt{2n}. \label{phi4}%
\end{equation}
Also,%
\[
\sqrt{2\pi}\left(  \frac{n}{a}\right)  ^{\frac{1}{6}}\left(  \sqrt{a}+\sqrt
{n}\right)  ^{\frac{1}{3}}\sim n^{\frac{1}{6}}%
\]
and%
\[
\left(  \frac{n}{a}\right)  ^{\frac{1}{6}}\frac{\left(  x-\Omega^{+}\right)
}{\left(  \sqrt{a}+\sqrt{n}\right)  ^{\frac{2}{3}}}\sim n^{\frac{1}{6}}%
\sqrt{2}\left(  \xi-\sqrt{2n}\right)  .
\]
Therefore,%
\begin{equation}
H_{n}(\xi)\sim\Lambda_{4}\left(  \xi\right)  \equiv\sqrt{2\pi}n^{\frac{1}{6}%
}\exp\left[  \Phi_{4}\left(  \xi\right)  \right]  \mathrm{Ai}\left[
n^{\frac{1}{6}}\sqrt{2}\left(  \xi-\sqrt{2n}\right)  \right]  . \label{h5}%
\end{equation}
Once again, we have
\[
\Lambda_{3}\left(  -\xi\right)  =(-1)^{n}\Lambda_{4}\left(  \xi\right)
\]

\item Region VI: $-\sqrt{2n}\ll\xi\ll\sqrt{2n}.$

From (\ref{c10}), we immediately obtain%
\[
H_{n}(\xi)\sim\Lambda_{5}\left(  \xi\right)  \equiv\Lambda_{1}\left(
\xi\right)  +\Lambda_{2}\left(  \xi\right)  .
\]
Since $-1<\frac{\xi}{\sqrt{2n}}<1,$ we set
\begin{equation}
\xi=\sqrt{2n}\sin\left(  \theta\right)  ,\quad-\frac{\pi}{2}<\theta<\frac{\pi
}{2}. \label{theta}%
\end{equation}
From (\ref{sigma}) we have%
\[
\sigma=\sqrt{2n}\cos\left(  \theta\right)  \mathrm{i.}%
\]
Thus,
\[
n\pi\mathrm{i}+\Phi_{1}\left(  \xi\right)  =\frac{n}{2}\left[  \ln\left(
2n\right)  -\cos\left(  2\theta\right)  \right]  +n\left[  \frac{1}{2}%
\sin\left(  2\theta\right)  +\theta-\frac{\pi}{2}\right]  \mathrm{i,}%
\]%
\[
\Phi_{2}\left(  \xi\right)  =\frac{n}{2}\left[  \ln\left(  2n\right)
-\cos\left(  2\theta\right)  \right]  -n\left[  \frac{1}{2}\sin\left(
2\theta\right)  +\theta-\frac{\pi}{2}\right]  \mathrm{i,}%
\]
and%
\[
U_{1}\left(  \xi\right)  =\frac{\exp\left(  \frac{\theta}{2}\mathrm{i}\right)
}{\sqrt{2\cos\left(  \theta\right)  }},\quad U_{2}\left(  \xi\right)
=\frac{\exp\left(  -\frac{\theta}{2}\mathrm{i}\right)  }{\sqrt{2\cos\left(
\theta\right)  }}.
\]
Hence,%
\begin{equation}
\Lambda_{5}\left[  \sqrt{2n}\sin\left(  \theta\right)  \right]  =\sqrt
{\frac{2}{\cos\left(  \theta\right)  }}\exp\left[  \Phi_{5}\left(
\theta\right)  \right]  \cos\left(  \Theta\right)  , \label{lambda5}%
\end{equation}
with%
\begin{equation}
\Phi_{5}\left(  \theta\right)  =\frac{n}{2}\left[  \ln\left(  2n\right)
-\cos\left(  2\theta\right)  \right]  \label{phi5}%
\end{equation}
and%
\begin{equation}
\Theta=n\left[  \frac{1}{2}\sin\left(  2\theta\right)  +\theta-\frac{\pi}%
{2}\right]  +\frac{\theta}{2}. \label{Theta}%
\end{equation}

Using (\ref{theta}), we can write (\ref{phi5}) and (\ref{Theta}) in terms of
$\xi$%
\[
\Phi_{5}\left(  \theta\right)  =\frac{n}{2}\left[  \ln\left(  2n\right)
-1\right]  +\frac{\xi^{2}}{2},
\]%
\[
\Theta=\frac{\xi}{2}\sqrt{2n-\xi^{2}}+\left(  n+\frac{1}{2}\right)
\arcsin\left(  \frac{\xi}{\sqrt{2n}}\right)  -n\frac{\pi}{2}.
\]
Also,%
\[
\sqrt{\frac{2}{\cos\left(  \theta\right)  }}=\sqrt{2}\left(  1-\frac{\xi^{2}%
}{2n}\right)  ^{-\frac{1}{4}}.
\]
Therefore,%
\begin{align}
\Lambda_{5}\left(  \xi\right)   &  =\sqrt{2}\left(  1-\frac{\xi^{2}}%
{2n}\right)  ^{-\frac{1}{4}}\exp\left\{  \frac{n}{2}\left[  \ln\left(
2n\right)  -1\right]  +\frac{\xi^{2}}{2}\right\} \label{lambda51}\\
&  \times\cos\left[  \frac{\xi}{2}\sqrt{2n-\xi^{2}}+\left(  n+\frac{1}%
{2}\right)  \arcsin\left(  \frac{\xi}{\sqrt{2n}}\right)  -n\frac{\pi}%
{2}\right]  .\nonumber
\end{align}
Considering the leading term of (\ref{lambda51}) as $n\rightarrow\infty,$ we
obtain%
\[
\Lambda_{5}\left(  \xi\right)  \sim\sqrt{2}\exp\left\{  \frac{n}{2}\left[
\ln\left(  2n\right)  -1\right]  +\frac{\xi^{2}}{2}\right\}  \cos\left(
n\frac{\pi}{2}-\xi\sqrt{2n}\right)
\]
in agreement with formula (4.14.9) in \cite{MR0350075}.
\end{enumerate}

\section{Zeros}

Let us denote by $\zeta_{1}^{n}>\zeta_{2}^{n}>\cdots>\zeta_{n}^{n}$ the zeros
of $H_{n}\left(  \xi\right)  ,$ enumerated in decreasing order. It then
follows from (\ref{lambda5}) that $\zeta_{j}^{n}=\sqrt{2n}\sin\left(  \tau
_{j}^{n}\right)  ,$ where $\tau_{j}^{n}$ satisfies%
\[
n\left[  \frac{1}{2}\sin\left(  2\tau_{j}^{n}\right)  +\tau_{j}^{n}-\frac{\pi
}{2}\right]  +\frac{\tau_{j}^{n}}{2}=\frac{\pi}{2}-j\pi,\quad1\leq j\leq n.
\]
We can rewrite the equation%
\[
n\left[  \frac{1}{2}\sin\left(  2t\right)  +t-\frac{\pi}{2}\right]  +\frac
{t}{2}=A
\]
as Kepler's equation%
\begin{equation}
E-\varepsilon\sin(E)=M,\label{Kepler}%
\end{equation}
with%
\begin{equation}
E=2t,\quad M=2\frac{2A+n\pi}{2n+1},\quad\varepsilon=-\frac{2n}{2n+1}.\label{M}%
\end{equation}
It is well known \cite{MR1349110} that the solution of (\ref{Kepler}) can be
expressed as a Kapteyn series%
\begin{equation}
E=M+2%
{\displaystyle\sum\limits_{k=1}^{\infty}}
\frac{1}{k}\mathrm{J}_{k}\left(  k\varepsilon\right)  \sin\left(  kM\right)
,\label{kapteyn}%
\end{equation}
where $\mathrm{J}_{k}\left(  \cdot\right)  $ is a Bessel function of the first kind.

Thus, using (\ref{M}) in (\ref{kapteyn}) with $A=\frac{\pi}{2}-j\pi,$ we
obtain%
\begin{equation}
\tau_{j}^{n}=\pi\frac{1+n-2j}{2n+1}+%
{\displaystyle\sum\limits_{k=1}^{\infty}}
\frac{1}{k}\mathrm{J}_{k}\left(  -\frac{2n}{2n+1}k\right)  \sin\left(
2\pi\frac{1+n-2j}{2n+1}k\right)  , \label{zeros}%
\end{equation}
for $1\leq j\leq n.$ Using the reflection formula \cite{atlas} $\mathrm{J}%
_{k}\left(  -x\right)  =\left(  -1\right)  ^{k}\mathrm{J}_{k}\left(  x\right)
,$ we can write (\ref{zeros}) as
\begin{equation}
\tau_{j}^{n}=\frac{\pi}{2}-\frac{\pi}{2}\left(  4j-1\right)  N^{-1}-%
{\displaystyle\sum\limits_{k=1}^{\infty}}
\frac{1}{k}\mathrm{J}_{k}\left[  \left(  1-N^{-1}\right)  k\right]
\sin\left(  \frac{4j-1}{N}k\pi\right)  , \label{zeros1}%
\end{equation}
where $N=2n+1.$




\end{document}